Applications of Differential Equation of Newton's law of cooling / warming

1
Expert's answer
2022-03-28T12:13:03-0400

The temperature of many objects can be modelled using a differential equation. Newton's law of cooling (or heating) states that the temperature of a body changes at a rate proportional to the difference in temperature between the body and its surroundings. It is a reasonably accurate approximation in some circumstances.


More precisely, let "T" denote the temperature of an object and "T_0" the ambient temperature. If "t" denotes time, then Newton's law states that:


"\dfrac{dT}{dt} =-k(T-T_0)"


where "k" is a positive constant. Thus, if the object is much hotter than its surroundings, then "T-T_0" is large and positive, so "\dfrac{dT} {dt}" is large and negative, so the object cools quickly. If the object is only slightly hotter than its surroundings, then "T-T_0" is small positive, and the object cools slowly. So a cup of hot coffee will cool more quickly if you put it in the refrigerator!